# Two-dimensional fluids via matrix hydrodynamics

### v3

@@ 6. A conjecture on the long-time behavior and integrability
The conjecture does not imply that steady state is reached (e.g. a minimal flow). Indeed, unsteadiness can persist indefinitely, without further mixing. From batches of numerical experiments [45], we conjecture that persistent unsteadiness in $\Omega_{+}(\omega_{0})$ is connected to integrable configuration of “vortex blob dynamics”, where the centers of mass of $n$ fixed-shape blobs interact (cf. Fig. 5(a)). Depending on the domain’s geometry, there is a threshold $n$ for which the dynamics is always integrable [46]. The underlying mechanism is then that vortex mixing continues until a large-scale almost integrable configuration of vortex blobs is reached. Being quasi-periodic, it then acts as a barrier for further mixing, trapping the vorticity in persistent unsteadiness. On $\mathbb{S}^{2}$, the thresholds are $n=3$ for generic initial data, and $n=4$ for vanishing angular momentum. Long-time numerical simulations with Zeitlin’s model align with this conjecture: for generic (smooth) initial data they typically settle on $n=3$ blobs (e.g. Fig. 1), and for vanishing momentum on $n=4$ blobs; see [45, Sec. 4] for details.
@@ Appendix A Proof of Theorem 4
Let $f\in L^{2}(\mathbb{S}^{2})$ and $g\in H^{3+\varepsilon}(\mathbb{S}^{2})$, we have that for any $\varepsilon>0$, $0<\alpha<2/3$, and $0<\beta<2/5$ there exist positive constants $C(\alpha)$, $C(\beta)$, $C(\alpha,\epsilon)$, and $C(\beta,\epsilon)$ such that
@@
has $\min\{N,2l_{1}-2,2l_{2}-2\}$ terms, which we repeatedly use in the following steps.
@@
The thesis then follow from the Grönwall lemma. ∎
@@ Appendix B Relation to Berezin-Toeplitz quantization
Since Dirac [15] outlined quantization theory, it has ramified into different mathematical developments. In this paper, we obtain quantization via unitary representation theory, which perhaps is the most direct route when applicable. Non-commutative geometry (cf. [12]) offers a broader viewpoint, and geometric quantization (cf.  [52]) yet another. In general, the various frameworks are non-equivalent and applicable under different conditions. But for the sphere they coincide; it is somehow the archetype of quantization (in non-commutative geometry the “fuzzy sphere”, and in geometric quantization the tautological complex line bundle over the Riemann sphere $\mathbb{C}P^{1}$, as we shall see).
@@
Let $M$ be a compact Kähler manifold. The objective is to construct a mapping (the Berezin-Toeplitz quantization) between the space of smooth functions $C^{\infty}(M,\mathbb{C})$ and a space of operators on a finite dimensional quantum state space (i.e., the Hilbert space of “wave-functions”). State space is constructed as holomorphic sections of a tensor power of a Hermitian line bundle over $M$. Whereas this construction appears abstruse at first, it is straightforward in the simplest setting $M=\mathbb{C}P^{1}$. As we shall see, this corresponds to quantization on $\mathbb{S}^{2}$ as presented in section 3.1.
@@
For $f\in C^{\infty}(\mathbb{S}^{2})$, let $T_{N}(f)$ denote the matrix quantization operator constructed in section 3.1 above. Then
$T_{N}(f)\simeq i\sqrt{\frac{N+1}{N-1}}\mathrm{BT}_{N-1}(f).$

### v4

@@ 6. A conjecture on the long-time behavior and integrability
The conjecture does not imply that steady state is reached (e.g., a minimal flow). Indeed, unsteadiness can persist indefinitely, without further mixing. From batches of numerical experiments [45], we conjecture that persistent unsteadiness in $\Omega_{+}(\omega_{0})$ is connected to integrable configuration of “vortex blob dynamics”, where the centers of mass of $n$ fixed-shape blobs interact (cf. Fig. 5(a)). Depending on the domain’s geometry, there is a threshold $n$ for which the dynamics is always integrable [46]. The underlying mechanism is then that vortex mixing continues until a large-scale almost integrable configuration of vortex blobs is reached. Being quasi-periodic, it then acts as a barrier for further mixing, trapping the vorticity in persistent unsteadiness. On $\mathbb{S}^{2}$, the thresholds are $n=3$ for generic initial data, and $n=4$ for vanishing angular momentum. Long-time numerical simulations with Zeitlin’s model align with this conjecture: for generic (smooth) initial data they typically settle on $n=3$ blobs (e.g. Fig. 1), and for vanishing momentum on $n=4$ blobs; see [45, Sec. 4] for details.
@@ Appendix A Proof of Theorem 4
Let $f\in H^{1+\epsilon}(\mathbb{S}^{2})$ and $g\in H^{3+\varepsilon}(\mathbb{S}^{2})$, we have that for any $\varepsilon>0$, $0<\alpha<2/3$, and $0<\beta<2/5$ there exist positive constants $C(\alpha)$, $C(\beta)$, $C(\alpha,\epsilon)$, and $C(\beta,\epsilon)$ such that
@@
has $\min\{N,2l_{1}-2,2l_{2}-2\}$ different values of $l_{3}$, which we repeatedly use in the following steps.
@@
The thesis then follows from the Grönwall lemma. ∎
@@ Appendix B Relation to Berezin-Toeplitz quantization
Since Dirac [15] outlined quantization theory it has ramified in different mathematical directions. In this paper, we obtain quantization via unitary representation theory, which perhaps is the most direct route when applicable. Non-commutative geometry (cf. [12]) offers a broader viewpoint, and geometric quantization (cf. [52]) yet another. The various frameworks are in general non-equivalent and applicable under different conditions. But for the sphere they coincide; it is somehow the archetype of quantization (in non-commutative geometry the “fuzzy sphere”, and in geometric quantization the tautological complex line bundle over the Riemann sphere $\mathbb{C}P^{1}$, as we shall see).
@@
Let $M$ be a compact Kähler manifold. The objective is to construct a mapping (the Berezin-Toeplitz quantization) between the space of smooth functions $C^{\infty}(M,\mathbb{C})$ and a space of operators on a finite dimensional quantum state space (i.e., the Hilbert space of “wave-functions”). State space is constructed as holomorphic sections of a tensor power of a Hermitian line bundle over $M$. The construction may appear abstruse at first, but is straightforward in the simplest setting $M=\mathbb{C}P^{1}$. As we shall see, this corresponds to quantization on $\mathbb{S}^{2}$ as presented in section 3.1.
@@
For $f\in C^{\infty}(\mathbb{S}^{2})$, let $T_{N}(f)$ denote the matrix quantization operator constructed in section 3.1 above. Then, under the identification $\mathbb{C}_{N-1}[z_{1},z_{2}]\simeq\mathbb{C}^{N}$,
$T_{N}(f)=i\sqrt{\frac{N+1}{N-1}}\mathrm{BT}_{N-1}(f).$